# The behavioral spillover effect: Modeling behavioral interdependencies in multi-pathogen dynamics

### v1

@@ 3.1 Numerical results
(b) Disease $B$ dominance
Figure 4: Persistence and dominance of diseases across one year. (a) Persistence of both diseases (purple) or only disease AA (red), which has the higher basic reproduction number. (b) Percentage of the first year that disease $B$ prevalence is above disease $A$ prevalence. For $\mathcal{R}_{0,A}=\mathcal{R}_{0,B}=3$ the diseases produce independent, identical outbreaks so neither are considered dominant. Dots correspond to combination of spillover ($s$) and basic reproduction number of disease $B$ ($\mathcal{R}_{0,B}$) used for plots in Figure 3. Note that for $\mathcal{R}_{0,A}=\mathcal{R}_{0,B}=3$, the strains are independent and identical, regardless of the level of spillover, so neither are considered dominant.
@@ 3.1.1 Approximation of spillover
To better understand the role of spillover in co-existence of both diseases, we assess our analytical conditions of stability equilibria (see Section 3.2 for their calculation) as well as provide an approximation of the minimum spillover required in order for disease $A$ to exclude disease $B$. See Theorem 1 below. The combination of the level of spillover and the basic reproduction number for disease $B$ (under our assumption that $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$) in order for co-existence of both diseases is seen in Figure 5. Compared to our numerical simulations in Figure 4a and our numerically-evaluated analytical conditions in Figure 5b, our approximation (seen in Figure 5a) under-estimates the parameter regime for which there is co-existence, i.e., co-existence is still possible for larger spillover and smaller basic reproduction number than the approximation suggests. Furthermore, there are slight differences between our simulation based results (Figure 4a) and our numerically-calculated analytical results (Figure 5b), such as for low $\mathcal{R}_{0,B}$ and low $s$, as we assess persistence of disease $B$ in the simulation results as falling below the threshold of $10^{-4}$ at one year.
@@
Figure 5: Approximated and analytical persistence of diseases. (a) Approximated threshold for co-existence of both diseases as determined in Theorem 1. (b) Analytically based computation of persistence of both diseases (purple) or only disease AA (red), which has the higher basic reproduction number. Dots correspond to combination of spillover ($s$) and basic reproduction number of disease $B$ ($\mathcal{R}_{0,B}$) used for plots in Figure 3.
@@
for exclusion of strain $B$.
To provide an intuition about the threshold, consider ${\mathcal{R}_{0,A}}=3$ and ${\mathcal{R}_{0,B}}=1.2$, which may resemble the case of SARS-CoV-2 and an influenza-like virus. In this case, $s_{threshold}=0.5$, such that for a higher values of spillover ($s>0.25$), there is complete suppression of disease $B$, the influenza-like virus (e.g., Figure 3, bottom row, middle column, where ${\mathcal{R}_{0,A}}=3$ and ${\mathcal{R}_{0,B}}=1.3$).
@@ 4 Discussion
This paper examines the dynamics of interdependencies across multiple pathogens in the absence of cross-immunity. While previous studies have often attributed inter-pathogen interactions to cross-immunity – where infection by one pathogen confers immunity against another [22, 23] – this study relaxes that assumption and instead highlights a behavioral mechanism that can couple the spread of two pathogens. This perspective is motivated by observations during the COVID-19 pandemic, where a surge in COVID-19 cases coincided with a marked decline in Influenza A and B cases [16, 17] This decline was largely due to public adherence to NPIs implemented to control COVID-19, which also inadvertently suppressed the spread of influenza [18]. We refer to this form of interdependence as behavioral spillover. This observation raises an important question: under what conditions can behavioral responses to one deadly disease contain the spread of others?
@@
To explore this, we develop a mathematical model to examine behavioral interdependencies across two pathogens. Building on a previously established behavioral epidemic model (SEIRSb) [35, 8], we extended the framework to capture the spread of two pathogens, where behavioral responses to one can influence the transmission dynamics of the other. Our model represents two diseases of different infectivity, and varying degrees of behavioral interdependence. Specifically, the model investigates the extent to which NPIs targeting one pathogen affect the spread of another.
@@
Our analysis explores a range of simulation outcomes for different values of the basic reproduction number and spillover strength. The results indicate that NPIs targeted at a pathogen with a relatively high basic reproduction number can significantly suppress the spread of other pathogens in the short term. This leads to scenarios in which one pathogen dominates while others maintain effective reproduction numbers below one, i.e., are fully suppressed. For example, in the presence of complete behavioral spillover, we observed that the population may only experience the disease with the highest $\mathcal{R}_{0}$, as if the ecosystem, including the diseases and humans, selects the pathogen with the highest $\mathcal{R}_{0}$. Additionally, we identify parameter regimes in which both pathogens coexist but spread at reduced rates compared to their behavior in isolation. With imperfect spillover, where NPIs designed for one disease only partially affect the spread of others, we estimate the range of values $\mathcal{R}_{0}$ in which only one of the pathogens spreads.
@@
Moreover, our study explored equilibria properties of the new model. Each scenario has four possible equilibria, depending on the values of the $\mathcal{R}_{0,i}$: one disease-free, two boundary (endemic in one disease and disease-free in the other), and one endemic equilibrium. While results regarding the disease-free equilibria were typical for disease models, i.e., the basic reproductive number functioned as a threshold value at one, results for the boundary and endemic equilibria when behavioral spillover was present diverged from this norm as one was not the threshold for stability for the equilibria. Within these scenarios, it is possible to have $\mathcal{R}_{0,A},\mathcal{R}_{0,B}>1$ with a stable boundary equilibrium and unstable endemic equilibrium; that is, the disease with the greater reproduction number will persist, while the other disease will die out despite both having a basic reproduction number above one, the typical threshold for persistence [4, 24, 52]. In cases with spillover, transmission depends on both diseases; analytically, the threshold condition for the newly introduced disease thus depends on the value of the infectious equilibrium of the endemic disease. If this threshold is large enough (i.e. if the presence of the endemic disease is large enough), then the newly introduced disease will be unsuccessful in establishing itself endemically. We identify the thresholds which determine the persistence disease(s) for all three scenarios.
@@
A critical component of working with a novel model is the parameterization with respect to real world data. This procedure usually involves a thorough literature review or fitting the model, which should not be done before conducting a structural identifiability analysis [50, 53]. Our identifiability analysis demonstrated that all the parameters are structurally locally identifiable with respect to prevalence and recognized prevalence data. The only scenario in which we obtained unidentifiable results was for the initial conditions of perceived prevalence, $\widetilde{I}_{A}(0)$ and $\widetilde{I}_{B}(0)$, when using prevalence data in the case of perfect spillover. Although $\widetilde{I}_{A}(0)$ and $\widetilde{I}_{B}(0)$ are unidentifiable, the initial conditions values are known, which provided confidence to proceed with the practical identifiability analysis. From the practical identifiability analysis, we observed that the majority of the parameters were identifiable in some form (i.e. strongly or weakly) with respect to prevalence and recognized prevalence. The immunity period, $\tau_{R}$, was the only parameter that was not identifiable. Since we used $\tau_{R}=100$ representing an immunity period of 100 days, and we simulated diseases dynamics for 100 days, it is probable that $\tau_{R}$ is identifiable if the simulations were lengthened considerably or if the immunity period was shortened.
@@
This study contributes to the fields of mathematical biology and epidemiological modeling, particularly to the emerging area of behavioral epidemic modeling [31]. Firstly, it adds to the body of literature on multi-pathogen dynamics [22, 23] by offering a contrasting perspective: that interdependent epidemic patterns can arise even in the absence of cross-immunity, purely due to behavioral interdependencies. This paper introduced the term behavioral spillover to describe this phenomenon, in which behavioral responses to one pathogen influence the transmission of others. An important implication is that ignoring such behavioral interdependencies may lead to misleading conclusions—such as over- or under-estimating pathogen transmission rate or wrongly inferring the presence of cross-immunity between pathogens. Second, this study contributes to endogenous formulation of human behavior [9]. It builds on prior behavioral epidemic models that incorporate a risk-response feedback loop, wherein rising (or falling) prevalence increases (or decreases) adherence to non-pharmaceutical interventions (NPIs), which in turn influences transmission [34, 39, 54]. The paper extends this framework to demonstrate how multiple pathogens can become dynamically coupled through shared behavioral responses.
@@
This study, however, has several limitations that point to directions for future research. First, the current model assumes a homogeneous population and does not account for heterogeneity in risk or behavioral response. For example, while elderly populations are at higher risk for both COVID-19 and Influenza A/B, other diseases like Lyme disease primarily affect different demographics, such as children. Incorporating population heterogeneity could yield more nuanced insights. Second, the model does not include vaccination dynamics, which can significantly alter behavioral interdependencies. For instance, after COVID-19 vaccines became available, increased perceived safety led to greater social activity, potentially raising the risk of influenza spread. Accounting for such vaccine-induced behavioral shifts would be a valuable extension.
@@
In conclusion, while the current model focuses on two pathogens for clarity, it can be generalized to consider multiple pathogens, different variants, and combinations of cross-immunity and behavioral spillover. Such extensions could offer broader insights into epidemic interactions in complex real-world settings. Furthermore, our focus was on offering a mathematical formulation of relevant models and the effort should be extended to include empirical tests using data across different regions and over a long period of time.

### v2

@@ 3.1 Numerical results
(b) Time disease $B$ exceeds disease AA
(c) Amount disease $A$ exceeds disease $B$
(d) Amount disease $B$ exceeds disease AA
Figure 4: Persistence and superiority of diseases across one year as computed through numerical simulation. (a) Persistence of both diseases (purple) or only disease AA (red), which has the higher basic reproduction number. (b) Percentage of the first year that disease $B$ prevalence is above disease AA prevalence. For $\mathcal{R}_{0,A}=\mathcal{R}_{0,B}=3$ the diseases produce independent, identical outbreaks so neither disease exceeds the other one (white line at the tope). White space at the bottom right corresponds to when there is persistence of only disease $A$. (c) Cumulative amount during the first year that disease $A$ exceeds disease $B$. (d) Cumulative amount during the first year that disease $B$ exceeds disease $A$. Note the different color scaling in panels (b)-(d). Dots correspond to combination of spillover ($s$) and basic reproduction number of disease $B$ ($\mathcal{R}_{0,B}$) used for plots in Figure 3.
@@ 3.1.1 Approximation of spillover
To better understand the role of spillover in co-existence of both diseases, we assess our analytical conditions of stability equilibria (see Section 3.2 for their calculation) as well as provide an approximation of the minimum spillover required in order for disease $A$ to exclude disease $B$. See Theorem 1 below. The combination of the level of spillover and the basic reproduction number for disease $B$ (under our assumption that $\mathcal{R}_{0,A}>\mathcal{R}_{0,B}>1$) in order for co-existence of both diseases is seen in Figure 5. Compared to our numerical simulations in Figure 4a and our numerically-evaluated analytical conditions in Figure 5b (which are visually nearly identical), our approximation (shown in Figure 5a) under-estimates the parameter regime for which there is co-existence, i.e., co-existence is still possible for larger spillover and smaller basic reproduction number than the approximation suggests. Furthermore, there are slight differences between our simulation based results (Figure 4a) and our numerically-calculated analytical results (Figure 5b), such as for low $\mathcal{R}_{0,B}$ and low $s$, as we assess persistence of disease $B$ in the simulation results as falling below the threshold of $10^{-4}$ at one year.
@@
Figure 5: Analytical approximation and analytical solution of regions of disease persistence. (a) Approximated threshold for co-existence of both diseases as determined in Theorem 1. (b) Analytical solution (numerically computed) of persistence of both diseases (purple) or only disease AA (red). Recall that disease AA has, by assumption, the higher basic reproduction number ($\mathcal{R}_{0,A}=3$). Dots correspond to combination of spillover ($s$) and basic reproduction number of disease $B$ ($\mathcal{R}_{0,B}$) used for plots in Figure 3.
@@
for the possibility of exclusion of strain $B$.
To provide an intuition about the threshold, consider ${\mathcal{R}_{0,A}}=3$ and ${\mathcal{R}_{0,B}}=1.3$ (as in the bottom row of Figure 3), which may resemble the case of SARS-CoV-2 and an influenza-like virus. In this case, $s_{threshold}\approx 0.35$, such that for higher values of spillover ($s>0.35$), there is the possibility of complete suppression of disease $B$, the influenza-like virus. We observe this suppression when $s=0.5$ in Figure 3, bottom row, middle column, where ${\mathcal{R}_{0,A}}=3$ and ${\mathcal{R}_{0,B}}=1.3$.
@@ 3.2 Equilibria and their stability
Even though epidemics exhibit complex dynamics and such systems may rarely remain at equilibrium due to external perturbations or seasonal forcing, equilibrium analysis provides insight into the underlying mechanisms of disease spread and helps guide intuition about the system’s dynamics.
@@ 4 Discussion
This paper examines the dynamics of interdependencies across multiple pathogens in the absence of cross-immunity. While previous studies have often attributed inter-pathogen interactions to cross-immunity – where infection by one pathogen confers immunity against another [6, 24] – this study relaxes that assumption and instead highlights a behavioral mechanism that can couple the spread of two pathogens. This perspective is motivated by observations during the COVID-19 pandemic, where a surge in COVID-19 cases coincided with a marked decline in Influenza A and B cases [59, 30]. This decline was largely due to public adherence to NPIs implemented to control COVID-19, which also inadvertently suppressed the spread of influenza [62]. This observation raises an important question: under what conditions can behavioral responses to one deadly disease contain the spread of others? We do not aim to reproduce the full complexity of epidemic dynamics here; rather, this work is theoretical investigation into behaviorally mediated interactions between diseases in a framework where analysis is possible.
@@
To this end, we develop a mathematical model to examine behavioral interdependencies across two pathogens. Building on a previously established behavioral epidemic model (SEIRb) [37, 57], we extend the framework to capture the spread of two pathogens, where behavioral responses to one can influence the transmission dynamics of the other. We refer to this form of interdependence as behavioral spillover. Our model represents two diseases of different infectivity, and varying degrees of behavioral interdependence. Specifically, the model investigates the extent to which NPIs targeting one pathogen affect the spread of another.
@@
Our analysis explores a range of simulation outcomes for different values of the basic reproduction number and spillover strength. The results indicate that NPIs targeted at a pathogen with a relatively high basic reproduction number can significantly suppress the spread of other pathogens in the short term. This leads to scenarios in which one pathogen exceeds the rest while others maintain effective reproduction numbers below one, i.e., are fully suppressed. For example, in the presence of complete behavioral spillover, we observed that the population may only experience the disease with the highest $\mathcal{R}_{0}$, as if the ecosystem, including the diseases and humans, selects the pathogen with the highest $\mathcal{R}_{0}$. Additionally, we identify parameter regimes in which both pathogens coexist but spread at reduced rates compared to their behavior in isolation. With imperfect spillover, where NPIs designed for one disease only partially affect the spread of others, we estimate the range of values $\mathcal{R}_{0}$ in which only one of the pathogens spreads.
@@
Moreover, our study explored equilibria properties of the new model given in System (LABEL:eqn:SIRS). Each scenario has four possible equilibria, depending on the values of the $\mathcal{R}_{0,i}$: one disease-free, two boundary (endemic in one disease and disease-free in the other), and one endemic equilibrium. While results regarding the disease-free equilibria were typical for disease models, i.e., the basic reproductive number functioned as a threshold value at one, results for the boundary and endemic equilibria when behavioral spillover was present diverged from this norm as one was not the threshold for stability for the equilibria. Within these scenarios, it is possible to have $\mathcal{R}_{0,A},\mathcal{R}_{0,B}>1$ with a stable boundary equilibrium and unstable endemic equilibrium; that is, the disease with the greater reproduction number will persist, while the other disease will die out despite both having a basic reproduction number above one, the typical threshold for persistence [7, 36, 26]. In cases with spillover, transmission depends on both diseases; analytically, the threshold condition for the newly introduced disease thus depends on the value of the infectious equilibrium of the endemic disease. If this threshold is large enough (i.e. if the presence of the endemic disease is large enough), then the newly introduced disease will be unsuccessful in establishing itself endemically. We identify the thresholds that determine the persistence disease(s) for all three scenarios.
@@
As a critical component of working with a novel model is the parameterization with respect to real world data. This procedure usually involves a thorough literature review or model fitting. To assess the potential utility of our novel model to be fit to data, we conducted a structural identifiability analysis [61, 67]. Our identifiability analysis demonstrated that all the parameters are structurally locally identifiable with respect to prevalence and recognized prevalence data. The only scenario in which we obtained unidentifiable results was for the initial conditions of perceived prevalence, $\widetilde{I}_{A}(0)$ and $\widetilde{I}_{B}(0)$, when using prevalence data in the case of perfect spillover. Although $\widetilde{I}_{A}(0)$ and $\widetilde{I}_{B}(0)$ are unidentifiable, the initial condition values are known, which provided confidence to proceed with the practical identifiability analysis. From the practical identifiability analysis, we observed that the majority of the parameters were identifiable in some form (i.e. strongly or weakly) with respect to prevalence and recognized prevalence. The immunity period, $\tau_{R}$, was the only parameter that was not identifiable. Since we used $\tau_{R}=100$ representing an immunity period of 100 days, and we simulated the diseases dynamics for 100 days, it is probable that $\tau_{R}$ is identifiable if the simulations were lengthened considerably or if the immunity period was shortened.
@@
This study contributes to the fields of mathematical biology and epidemiological modeling, particularly to the emerging area of behavioral epidemic modeling [21]. Firstly, it adds to the body of literature on multi-pathogen dynamics [6, 24] by offering a contrasting perspective: that interdependent epidemic patterns can arise even in the absence of cross-immunity, purely due to behavioral interdependencies. This paper introduced the term behavioral spillover to describe this phenomenon, in which behavioral responses to one pathogen influence the transmission of others. An important implication is that ignoring such behavioral interdependencies may lead to misleading conclusions—such as over- or under-estimating pathogen transmission rate or wrongly inferring the presence of cross-immunity between pathogens. Second, this study contributes to the endogenous formulation of human behavior [38]. It builds on prior behavioral epidemic models that incorporate a risk-response feedback loop, wherein rising (or falling) prevalence increases (or decreases) adherence to non-pharmaceutical interventions (NPIs), which in turn influences transmission [56, 20, 42]. The paper extends this framework to demonstrate how multiple pathogens can become dynamically coupled through shared behavioral responses.
@@
However, this study has several limitations that point to directions for future research. First, the current model assumes a homogeneous population and does not account for heterogeneity in risk or behavioral response. For example, while elderly populations are at higher risk for both COVID-19 and Influenza A/B, other diseases like Lyme disease primarily affect different demographics, such as children. Incorporating population heterogeneity could yield more nuanced insights. Second, the model does not include vaccination dynamics, which can significantly alter behavioral interdependencies. For instance, after COVID-19 vaccines became available, increased perceived safety led to greater social activity, potentially raising the risk of influenza spread. Accounting for such vaccine-induced behavioral shifts would be a valuable extension. Finally, the model assumes infected individuals do not alter their behavior and does not include biological cross-immunity between pathogens. While these assumptions simplify the system, they allow us to isolate behavioral mechanisms, showing that apparent protection against one disease can arise purely from changes in perceived risk of the other. Relaxing these assumptions—by incorporating quarantine upon infection or partial cross-immunity—could further strengthen the interdependencies and accelerate the practical extinction of diseases with lower reproduction numbers.
@@
In conclusion, while the current model focuses on two pathogens for clarity, it can be generalized to consider multiple specific pathogens, different variants, and combinations of cross-immunity and behavioral spillover. Such extensions could offer broader insights into epidemic interactions in complex real-world settings. Furthermore, our focus was on offering a mathematical formulation of relevant models and the effort should be extended to include empirical tests using data across different regions and over a long period of time.